OpenStudy (anonymous):

Simplify. a^5(a^4) = a_________

OpenStudy (anonymous):

@mathmale

OpenStudy (esshotwired):

When you multiply exponents, you add the exponents together, but the number that's not an exponent stays the same. Try figuring it out now

OpenStudy (mathmale):

@esshotwired: Please help Starr_DynastyT to figure it out himself / herself.

OpenStudy (esshotwired):

I just told them how to do it on their own, and if they need, more help they will ask.

OpenStudy (anonymous):

\[a^5(a^4) = a^{5+4} = a^90\] Since the base are same ..the exponents will add up

OpenStudy (anonymous):

there's a tyo.. answer is a ^9

OpenStudy (anonymous):

So i added it and it got me the answer 9

OpenStudy (anonymous):

is that correct?..if yes from now on try to do this kind of question on yar' own

OpenStudy (anonymous):

i rather it explained but thanks anyway @Adjax

OpenStudy (texaschic101):

exponent rule : a^m * a^n = a ^ (m + n)

OpenStudy (mathmale):

"So i added it and it got me the answer 9" should read, "I added up the two exponents and obtained 9, so my final result is a^9."